a merchant has coffee worth $20 a pound that she wishes to mix with 90 pounds of coffee worth $60 a pound to get a mixture that can be sold for $30 a pound. How many pounds of the $20 coffee should be used?
Answer by mananth(16946)

coffee $20 a pound x lbs
90 pounds of coffee worth $60 a pound
mixture x+90 lbs $30 a pound.
...
20x+90*60=30(x+90)
20x+5400=30x+2700
30x-20x=5400-2700
10x=2700
/10
x= 2700/10
x= 270 lbs of $20